OpenStudy (anonymous):

how you evaluate the expression 42 + 5 × 32 – 42 ÷ 23

OpenStudy (larseighner):

First do everything in parenthesis (you have none). Then do exponents (you have none). Then multiplications (and division). Finally addition (and subtraction).

OpenStudy (arnavguddu):

try 5*32=a, 42/23=b.... next 42+a=c next c-b=ur answr

OpenStudy (chrisfso):

so following Lars would be 42+(5x32)-(42/23) evaluate the contents of the parenthesis first, and then sum the three remaining terms. = 42+160-1.826 =200.174